How do I get the answer to 2y^2-5y-12
Found 2 solutions by oberobic, chessace:
Answer by oberobic(2304)

The expression can be factored:
(2y +3)(y -4)
.
So y = -3/2 or 4

Answer by chessace(471)

Factor into (2y + a)(y - b)
then a*b = 12, 2b-a = 5
b=4 a=3
If the question is "factor", Ans: (2y+3)(y-4)
If the question is "find roots", Ans: y = -3/2, 4
Next time give the question.